# The Discovery of Little Red Dots in the Local Universe: Signatures of Cool Gas Envelopes

### v1

@@ 8 Summary
We searched the SDSS spectroscopic database for local LRDs using high-redshift LRD templates and identified three LRDs at –. We followed up these sources with MMT/Binospec, LBT/MODS, and Magellan/FIRE spectroscopy, which reveal detailed spectral features that probe the physical nature of these local LRDs. Our findings are summarized below.
@@
The three local objects closely resemble the high-redshift LRDs. They display compact morphologies. HST narrow-band imaging of J1047+0739 reveals that both the continuum and the H line emission are AGN-dominated, with the compact yet clumpy host galaxy (half-light radius about 530 pc) contributing a minor fraction to the continuum. All the three objects exhibit clear V-shaped UV-to-optical continua with inflections around the Balmer break, and a decreasing IR continuum from Å to 2 µm. All of them show broad H emission with FWHMs km s-1, along with broad Paschen and He I lines. All of them exhibit blueshifted He I 10830 absorbers; two of them show H absorbers, and one even shows a redshifted H absorber without an H counterpart. None of the objects shows significant optical or IR variability. Two were targeted in hard X-rays but were either not detected or detected with low statistical significance.
@@
The Balmer and Paschen decrements, covering from H to H12 and from Pa to Pa12 (or Pa9 for one object), indicate negligible dust attenuation in the NLRs. The broad H/H ratios are in all three objects, higher than those observed in most local AGNs, indicating either unusually high-density gas or atypical dust properties in the BLRs.
@@
The BPT diagram shows that the narrow emission lines are photoionized by AGNs and overlap with the region occupied by a known LRD. The NLRs are metal-poor, with 12+(O/H)7.43–8.00.
@@
Two of the objects exhibit prominent narrow [Fe II] emission lines that are distinct from those typically seen in type-1 AGNs. Photoionization modeling suggests that the [Fe II]-emitting gas is AGN-powered, with physical conditions characterized by , , , and an average temperature of K. Based on the inferred gas density and FWHMs, we propose that the [Fe II]-emitting region lies at an intermediate scale between the BLRs and NLRs.
@@
One of the local LRDs, J1025+1402 (The Egg) exhibits exceptionally high EWs in Na D, K I, Fe II, and Ca T absorption lines, with narrow FWHMs of about 300–400 km s-1 and no significant velocity shifts. Such strong metal absorbers are rarely seen in the ISM, galaxies, AGNs, or stars. Its Ca T EW is matched only by that of two supergiant and hypergiant stars. These extreme properties suggest a non-stellar and non-ISM origin. We interpret the Ca T absorption in The Egg as arising from a cool gas envelope around the BH, analogous to stellar atmospheres. We infer – and at the photosphere, corresponding to a density of g cm-3.
@@
The optical-IR continua of these local LRDs exhibit pseudo-blackbody shapes. The shapes are consistent with theoretical models featuring atmospheres surrounding BHs (H. Liu et al., 2025), with characteristic temperatures of K. The WISE photometry indicates modest dust emission at temperatures of – K.
@@
By combining the observational evidence, we propose a conceptual model for the structure of LRDs. In this scenario, BHs are embedded within a cool gas envelope that emits thermalized radiation, giving rise to the observed pseudo-blackbody continuum and metal absorption features. The Balmer and He I absorbers are attributed to optically thick outflows launched from the envelope. The overall AGN structure consists of BLRs, an intermediate region that contains [Fe II]-emitting gas, and NLRs with minimal dust attenuation. The geometry of these structures, along with the viewing angle, may account for the diverse observed features across different objects.
@@
Local LRDs offer a valuable opportunity to investigate the physical nature of this population. The conceptual framework presented in §7 provides one possible scenario to explain the observed features. However, open questions remain. Do LRDs, and the scenario illustrated in Figure 14, represent a specific evolutionary phase of SMBHs? Are they connected to conventional type-1 and type-2 AGNs? How do the physical conditions that give rise to LRDs evolve with redshift, and why are they so rare in the local Universe? How do the BH and host galaxy properties of LRDs change over cosmic time? While answering these questions is beyond the scope of this paper, they are critical for understanding the broader context of BHs and merit future investigation.
@@
Ongoing wide-field surveys, such as DESI and Euclid, along with upcoming missions like the Roman Space Telescope, will open new windows for identifying LRDs at low redshift. These efforts will enable studies of their cosmic evolution and contributions to the history of SMBH growth. Large statistical samples will be essential for establishing their demographics. Future observations with high-resolution facilities such as HST and JWST will provide critical insights to test the proposed scenarios and directly probe the properties of their host galaxies.
@@ Appendix A HST FR782N imaging of J1047+0739
We model the morphology of the continuum map using galfit (C. Y. Peng et al., 2002). We assume a PSF and four Sérsic components. The right panel of Figure A.1 presents the best-fit model and the PSF-subtracted components, which represent the host galaxy. The host galaxy has compact, yet clumpy morphology, with a half-light radius of 530 pc. The luminosity of the central PSF is 5.5 times that of the host galaxy, indicating that the AGN contributes about 85% of the total continuum light in the optical.
@@ Appendix B Emission line fitting
For strong emission lines with visible broad components (e.g., H, Pa), we model each line with a single narrow Gaussian plus one or more broad Gaussian or Lorentzian profiles to capture the composite broad features. A linear polynomial is added to account for the local continuum. Specifically, broad Paschen lines are modeled with Lorentzian profiles, leveraging the high-resolution FIRE spectra to resolve line wings, while broad H components are depicted by two to three Gaussian profiles. The [N II] 6548, 6583 doublet is fit simultaneously with H. We model them with two narrow Gaussians, fixing their FWHMs to match the narrow H component and setting their flux ratio to 3. In J1047+0739, the Paschen lines and adjacent IR He I lines are modeled simultaneously, with the He I emission represented by broad and narrow Gaussian components. Note that for each line profile fit, the parameters of the narrow Gaussian component are treated independently, as the spectral resolution varies across wavelengths and instruments. For example, [O III]5007 in the MODS spectra is unresolved, with its observed FWHM determined by the wavelength-dependent line spread function. Therefore, we do not use its fit to constrain the narrow components of H or Pa.
For the H emission lines of The Egg and J1022+0841, whether from MODS or Binospec spectra, we simultaneously fit the absorption and emission profiles. Given the higher resolution of the Binospec spectra, we rely on them for more precise absorber measurements. As shown in Figure B.1, a Gaussian absorption component (EW Å) is applied on top of the composite broad H profile (i.e., the sum of the two broad Gaussians). For J1022+0841, we test two different fitting models: one assuming a single Gaussian absorber superimposed on the broad emission lines, and the other using two Gaussian absorbers. As shown in Figure B.1, the two-Gaussian model provides a better fit, more accurately capturing the profile wing.
Since the He I 10830 absorption profiles are complex, we model the intrinsic emission by masking the absorption-affected regions. The remaining line profile wings, unaffected by absorption, are used to constrain and recover the symmetric emission line profiles. The emission is modeled with one or two broad Gaussian components plus a narrow component. For H line of The Egg, we also mask the absorbed region and fit the intrinsic emission profile.
@@
The normalized H absorption profiles shown in Figure 6 are derived by dividing the absorbed broad H emission by the intrinsic broad H line profile. For H in The Egg and He I 10830 in all three objects, the profiles are normalized by the total unabsorbed profiles, including both the broad emission lines and continua. Normalizing H using only the broad emission line is tested but results in an absorption peak below zero. We then model the normalized H absorber with a single Gaussian and the normalized He I 10830 absorber with multiple Gaussians. Table 4 summarizes the best-fit parameters for each absorber.
Name | H | H | He I 10830 | ||||||
FWHM | EW | FWHM | EW | FWHM | EW | ||||
The Egg | 73 | ||||||||
(J1025+1402) | |||||||||
total | |||||||||
J1047+0739 | – | – | – | – | – | – | |||
total | |||||||||
J1022+0841 | – | – | – | (two Gaussian) | |||||
total | |||||||||
(one Gaussian) | |||||||||
@@
To model AGN photoionization scenarios for [Fe II] emission, we adopt AGN SEDs from C. Jin et al. (2012). While the chosen SED corresponds to an Eddington ratio of 0.17, our cloudy modeling is intended to be qualitative and is not highly sensitive to this assumed value. How the Eddington ratio impacts the AGN SED, and consequently the resulting [Fe II] spectrum, is beyond the scope of this paper. Our cloudy models span a grid of hydrogen densities from to (0.5 dex step), column densities from to (0.5 dex step), ionization parameters from to (1 dex step), and metallicities from to (1 dex step). We run a suite of cloudy models over this parameter space, assuming a plane-parallel geometry, and obtain the predicted [Fe II] line spectra.
To determine the best-fit models, we compare the observed emission line fluxes with those predicted by the modeled [Fe II] spectrum. Specifically, we use the fluxes of [Fe II] 4235, 4244, 4278, 4287, 4416, 5158, 5262, 7155, and 7172. Additionally, since [Fe II] 5169 is highly sensitive to the gas conditions but is not detected, we include its 3 upper limit as an extra constraint. All line fluxes, both observed and modeled, are normalized to that of [Fe II] 4244, and we compute the value for each model grid.
For comparison, we also explored models using a 5000 K blackbody as the ionizing source. The temperature is chosen based on our analysis in §6 and §7. These models fail to reproduce the observed [Fe II] line ratios, with significant discrepancies particularly in [Fe II] 4235, 4244, 4278, 4287, and 4416.
@@ D.1 Theoretical atmosphere model + dust templates
We begin with the best-matched theoretical models from H. Liu et al. (2025) and model the dust emission using the dust templates from J. Lyu & G. H. Rieke (2021). These dust templates, originally developed for type-1 AGNs, include four components: blackbodies at 2100 K, 1038 K, and 300 K, and a graybody at 90 K. We extrapolate the theoretical model spectra by assuming a blackbody tail in the near-IR, consistent with the model’s effective temperature. We then fit the amplitudes of the four dust components to reproduce the overall continuum and WISE photometry at wavelengths longer than 1.5 µm.
Figure 13 demonstrates that this model successfully reproduces the near-infrared photometry of The Egg. Figure D.1 shows the SEDs of J1047+0739 and J1022+0841 compared with and the models from H. Liu et al. (2025) that visually match the data. The intrinsic luminosity of the H. Liu et al. (2025) model is erg s-1 for J1047+0739, with a reference density of g cm-3 (which is comparable to the photosphere density; see H. Liu et al. (2025) for exact definition). These model parameters give an effective temperature of K. For J1022+0841, the model luminosity is erg s-1 with g cm-3, and the resulting K. We note that the SED of J1047+0739 is less well reproduced, mainly due to the non-negligible galaxy light. The HST narrow-band imaging (see Appendix A) reveals that, although the AGN dominates the continuum, the host galaxy contributes at least 18% of the continuum light around the H wavelengths. The contribution from the host galaxy is likely minor in The Egg and J1022+0841, both of which exhibit more prominent V-shaped SEDs with redder optical continua. This suggests stronger AGN dominance in the optical continua. Nevertheless, the comparison between the model and observed spectra aims to demonstrate the plausibility of the theoretical models rather than to constrain the physical properties of these objects.
@@ D.2 An empirical parameterization of the continuum
We attempt to empirically parameterize the optical-IR continuum shapes of local LRDs with a pseudo-blackbody spectrum. This parameterization is purely descriptive, intended to capture the observed continuum shape but not to provide a physical interpretation. We adopt the functional form of a modified Calzetti law (S. Noll et al., 2009), but it should not be interpreted as indicating any particular amount of dust attenuation. The parameterization can be expressed as:
where the free parameters are , , , , , , and . and are the amplitude and power-law slope of the UV continuum. It represents a distinct component, which could be either due to the AGN light or the host galaxy. denotes the form of the modified Calzetti law (S. Noll et al., 2009), where represents a generalized attenuation curve (not necessarily tied to physical dust properties), and is the power-law modification parameter, normalized at 5500Å. For J1047+0739, we introduce an additional free parameter to model the Balmer jump as a discontinuity in the power-law at 3645 Å. is the blackbody spectrum at temperature , and is its amplitude. For the dust emission, we adopt the dust templates from J. Lyu & G. H. Rieke (2021). and are the amplitude and blackbody or gray body spectrum of the -th dust component.
The best-fit continua of the three local LRDs are shown in Figure D.2. The corresponding best-fit parameters and component luminosities are summarized in Table 5. We do not include associated uncertainties given the descriptive nature of this parameterization. All three objects exhibit temperatures around 5000–6000 K and very negative values of . The extreme produces a very steep modification to the original blackbody. Such a negative strongly suppresses the UV-to-optical flux of the blackbody spectrum while leaving the IR spectrum largely unaffected. As discussed in §6.2, physical mechanisms, such as metal cooling and gas opacity, may contribute to this modification. In particular, the contribution of the UV power-law component to the optical continuum in J1047+0739 is non-negligible. At the H wavelength, the power-law component is only a factor of 2.3 fainter than the pseudo-blackbody. This is broadly consistent with measurements from its HST narrow-band image, which show the AGN luminosity to be 5.5 times that of the host galaxy. In the other two objects, the power-law components at optical wavelengths are orders of magnitude fainter than the pseudo-blackbody components.

### v2

@@ 8 Summary
We searched the SDSS spectroscopic database for local LRDs using high-redshift LRD templates and identified three LRDs at –. Their discovery suggests a conservative lower limit for LRD number density at of . We followed up these sources with MMT/Binospec, LBT/MODS, and Magellan/FIRE spectroscopy, which reveal detailed spectral features that probe the physical nature of these local LRDs. Our findings are summarized below.
@@
The three local objects closely resemble the high-redshift LRDs. They display compact morphologies (0.85-1.6 pkpc). HST narrow-band imaging of J1047+0739 reveals that both the continuum and the H line emission are AGN-dominated, with the compact yet clumpy host galaxy ( pc) contributing a minor fraction to the continuum. All three objects exhibit clear V-shaped UV-to-optical continua with inflections near the Balmer break, and a decreasing IR continuum from Å to 2 µm. All of them show broad H emission with FWHMs km s-1, along with broad Paschen and He I lines. All of them exhibit blueshifted He I 10833 absorbers; two of them show H absorbers, and one even shows a redshifted H absorber without an H counterpart. None of the objects shows significant optical or IR variability. Two were targeted in hard X-rays but were either not detected or detected with low statistical significance.
@@
The Balmer and Paschen decrements, covering from H to H12 and from Pa to Pa12 (or Pa9 for one object), indicate negligible dust attenuation in the NLRs, consistent with Case B recombination. The broad H/H ratios are 8-34 in the three objects, higher than those observed in most local AGNs, indicating either unusually high-density gas or atypical dust properties in the BLRs.
@@
The BPT diagram shows that the narrow emission lines are photoionized by AGNs, with the three LRDs occupying positions consistent with those reported by Juodzbalis2024. Their NLRs are metal-poor, with 12 + (O/H) 7.43–8.01.
@@
Two of the objects exhibit prominent narrow [Fe II] (or Fe II) emission lines that are distinct from those typically seen in type-1 AGNs. Photoionization modeling suggests that the [Fe II]-emitting gas is AGN-powered, with physical conditions characterized by , , , and an average temperature of K. Based on the inferred gas density and FWHMs, we propose that the [Fe II]-emitting region lies at an intermediate scale between the BLRs and NLRs.
@@
One of the local LRDs, J1025+1402 (The Egg) exhibits exceptionally high EWs in Na D, K I, and Ca T absorption lines, with FWHMs of about 300–440 km s-1 and no significant velocity shifts. Such strong metal absorbers are rarely seen in the ISM, galaxies, AGNs, or stars. Its Ca T EW is matched only by that of two supergiant and hypergiant stars. The Egg also exhibits potential low-ionization absorbers, such as Fe I, Ti II, Mg I, Ba II, etc., characteristic of G- to K-type stars. These extreme properties suggest a non-stellar and non-ISM origin. We interpret these absorption features in The Egg as arising from a cool gas envelope around the BH, analogous to stellar atmospheres. Based on its Ca T EW, we infer – and at the photosphere, corresponding to a density of g cm-3.
@@
The optical-IR continua of these local LRDs exhibit pseudo-blackbody shapes. The shapes are consistent with theoretical models featuring atmospheres surrounding BHs (Liu2025), with characteristic temperatures of K. The WISE photometry indicates modest dust emission at temperatures of – K.
@@
By combining the observational evidence, we propose a conceptual model for the structure of LRDs. In this scenario, BHs are embedded within a cool gas envelope that emits thermalized radiation, giving rise to the observed pseudo-blackbody continuum and metal absorption features. The Balmer and He I absorbers are attributed to optically thick outflows launched from the envelope. The overall AGN structure consists of BLRs, an intermediate region that contains [Fe II]-emitting gas, and NLRs with minimal dust attenuation. The geometry of these structures, along with the viewing angle, may account for the diverse observed features across different objects (i.e., the presence or absence of cool star–like absorption).
@@
Local LRDs offer a valuable opportunity to investigate the physical nature of this population. The conceptual framework presented in §7 provides one possible scenario to explain the observed features. However, open questions remain. Do LRDs, and the scenario illustrated in Figure 15, represent a specific evolutionary phase of SMBHs? What is the accretion mode of these objects (super-Eddington or sub-Eddington)? Are they connected to conventional type-1 and type-2 AGNs? How do the physical conditions that give rise to LRDs evolve with redshift, and why are they so rare in the local Universe? How do the BH and host galaxy properties of LRDs change over cosmic time? While answering these questions is beyond the scope of this paper, they are critical for understanding the broader context of BHs and merit future investigation.
@@
Ongoing wide-field surveys, such as DESI and Euclid, along with upcoming missions such as the Prime Focus Spectrograph (PFS), LSST, and Roman Space Telescope, will open new windows for identifying LRDs at low redshift. These efforts will enable studies of their cosmic evolution and contributions to the history of SMBH growth. Large statistical samples will be essential for establishing their demographics. Future observations with high-resolution facilities such as HST and JWST will provide critical insights to test the proposed scenarios and directly probe the properties of their host galaxies.
@@ Appendix A HST FR782N imaging of J1047+0739
We model the morphology of the continuum map using galfit (Peng2002). We assume a PSF and four Sérsic components, with the number of Sérsic components determined through several trials; four Sérsic yielded the optimal fit. The right panel of Figure A.1 presents the best-fit model and the PSF-subtracted components, which represent the host galaxy. The host galaxy has compact, yet clumpy morphology, with a half-light radius of 568 pc. The luminosity of the central PSF is 5.5 times that of the host galaxy, indicating that the AGN contributes about 85% of the total continuum light in the optical.
@@ Appendix B Emission line fitting
For the H emission lines of The Egg and J1022+0841, we use observations from LBT/MODS and MMT/Binospec. Given the higher resolution of Binospec (–5300, compared to –1300 for MODS), we adopt the Binospec spectra for the fiducial fits and degrade the best-fit models to the MODS resolution for consistency checks. In the Binospec spectra, the absorption and emission components are fitted simultaneously, with a linear polynomial included to account for the local continuum. The [N II] 6550, 6585 doublet is modeled jointly with H using two narrow Gaussians, with their FWHMs tied to that of the narrow H component and their flux ratio fixed to 3. As shown in Figure B.1, the best-fit models reproduce the Binospec spectra well. When degraded to the resolution of MODS, the models remain in good agreement with the MODS H profiles. Small discrepancies are present in the broad wings of The Egg, arising from the fact that the instrumental line-spread functions might not be perfectly Gaussian. We verify that such discrepancies introduce only a 5% difference in the broad H luminosity when fitting the MODS spectrum alone, and thus have a negligible impact on the derived physical properties (e.g., ). For J1022+0841, we test two different fitting models: one assuming a single Gaussian absorber superimposed on the broad emission lines, and the other using two Gaussian absorbers. As shown in Figure B.1, the two-Gaussian model provides a better fit, more accurately capturing the wing of the absorption profile. The fitting results using the two-Gaussian model are reported in Table 3. For the single-Gaussian model, we obtain km s-1, FWHM= km s-1, and EW= Å. The normalized H absorption profiles shown in Figure 6 are derived by dividing the absorbed broad H emission by the intrinsic broad H line profile.
@@
For strong emission lines exhibiting visible broad components (e.g., H, Pa) in the LBT/MODS and Magellan/FIRE spectra, each line is modeled with a single narrow Gaussian plus one or more broad Gaussian or Lorentzian profiles to capture the composite broad features. A linear polynomial is added to account for the local continuum. Specifically, broad Paschen lines are modeled with Lorentzian profiles, leveraging the high-resolution FIRE spectra to resolve line wings. In J1047+0739, the Paschen lines and adjacent IR He I lines are modeled simultaneously, with the He I emission represented by broad and narrow Gaussian components. Note that for each line profile fit, the parameters of the narrow Gaussian component are treated independently, as the spectral resolution varies across wavelengths and instruments. For example, [O III] 5008 in the MODS spectra is unresolved, with its observed FWHM determined by the wavelength-dependent line spread function. Therefore, we do not use its fit to constrain the narrow components of H or Pa. The measurements of key emission lines are summarized in Table B.1.
| J1025+1402 | J1047+0739 | J1022+0841 | |
| ( erg s-1) | |||
| (km s-1) | – | – | |
| ( erg s-1) | |||
| ( erg s-1) | |||
| ( erg s-1) | |||
| ( erg s-1) | |||
| FWHMHβ,broad (km s-1) | |||
| ( erg s-1) | |||
| FWHMPaα,broad (km s-1) | |||
| ( erg s-1) | |||
| FWHMPaα,narrow (km s-1) | |||
| ( erg s-1) | |||
| FWHMPaβ,broad (km s-1) | a | ||
| ( erg s-1) | |||
| FWHMPaβ,narrow (km s-1) | – a | ||
| ( erg s-1) | |||
| FWHMPaγ,broad (km s-1) | |||
| ( erg s-1) | |||
| FWHMPaγ,narrow (km s-1) | |||
| ( erg s-1) | |||
| FWHMPaδ,broad (km s-1) | |||
| ( erg s-1) | |||
| FWHMPaδ,narrow (km s-1) | – |
@@
For H line of The Egg, we mask the absorbed region and fit the intrinsic emission profile. Similarly, the intrinsic He I 10833 emission is modeled by excluding regions affected by absorption. The unaffected line wings are used to constrain and recover the symmetric emission profile. The emission is modeled with one or two broad Gaussian components plus a narrow component. In Figure 6, the H profile in The Egg and the He I 10833 profiles in all three objects are normalized by the total unabsorbed profiles, including both the broad emission lines and continua. Normalizing H using only the broad emission line is tested, but results in an absorption peak below zero. We then model the normalized H absorber with a single Gaussian and the normalized He I 10833 absorber with multiple Gaussians. Table 3 summarizes the best-fit parameters for each absorber.
@@
To model AGN photoionization scenarios for [Fe II] emission, we adopt AGN SEDs from Jin2012. While the chosen SED corresponds to an Eddington ratio of 0.17, our cloudy modeling is intended to be qualitative and is not highly sensitive to this assumed value. How the Eddington ratio impacts the AGN SED, and consequently the resulting [Fe II] spectrum, is beyond the scope of this paper. Our cloudy models span a grid of hydrogen densities from to (0.5 dex step), column densities from to (0.5 dex step), ionization parameters from to (1 dex step), and metallicities from to (1 dex step). We run a suite of cloudy models over this parameter space, assuming a plane-parallel geometry, and obtain the predicted [Fe II] line spectra.
To determine the best-fit models, we compare the observed emission line fluxes with those predicted by the modeled [Fe II] spectrum. Specifically, we use the fluxes of Fe II 4234, [Fe II] 4245, 4278, 4289, 4416 (blend of 4415 and 4418), 5159, 5263, 7157, and 7174. Additionally, since Fe II 5171 is highly sensitive to the gas conditions but is not detected, we include its 3 upper limit as an extra constraint. All line fluxes, both observed and modeled, are normalized to that of [Fe II] 4245, and we compute the value for each model grid.
For comparison, we also explored models using a 5000 K blackbody as the ionizing source. The temperature is chosen based on our analysis in §6 and §7. These models fail to reproduce the observed [Fe II] line ratios, with significant discrepancies particularly in Fe II 4234, [Fe II]4245, 4278, 4289, and 4416.
@@ D.1 Theoretical atmosphere model + dust templates
We begin with the best-matched theoretical models from Liu2025 and model the dust emission using the dust templates from Lyu2021. As demonstrated in Lyu2017 and Lyu2021, typical AGN-heated dust emission SED can be well can be well characterized by the combination of four major dust components: blackbodies at 2100 K, 1038 K, and 300 K, and a graybody at 90 K. We extrapolate the theoretical model spectra by assuming a blackbody tail in the near-IR, consistent with the model’s effective temperature. We then fit the amplitudes of the four dust components to reproduce the overall continuum and WISE photometry at wavelengths longer than 1.5 µm.
Figure 14 demonstrates that this model successfully reproduces the near-infrared photometry of The Egg. Figure D.1 shows the SEDs of J1047+0739 and J1022+0841 compared with and the models from Liu2025 that visually match the data. The intrinsic luminosity of the Liu2025 model is erg s-1 for J1047+0739, with a reference density of g cm-3 (which is comparable to the photosphere density; see Liu2025 for exact definition). These model parameters give an effective temperature of K. For J1022+0841, the model luminosity is erg s-1 with g cm-3, and the resulting K. We note that the SED of J1047+0739 is less well reproduced, possibly due to the non-negligible galaxy light. Indeed, the HST narrow-band imaging (see Appendix A) reveals that, although the AGN dominates the continuum, the host galaxy contributes about 15% of the continuum light around the H wavelengths. The contribution from the host galaxy is likely minor in The Egg and J1022+0841, both of which exhibit more prominent V-shaped SEDs with redder optical continua. This suggests stronger AGN dominance in the optical continua. Nevertheless, the comparison between the model and observed spectra aims to demonstrate the plausibility of the theoretical models rather than to constrain the physical properties of these objects.
@@ D.2 An empirical parameterization of the continuum
We attempt to empirically parameterize the optical-IR continuum shapes of local LRDs with a pseudo-blackbody spectrum. This parameterization is purely descriptive, intended to capture the observed continuum shape but not to provide a physical interpretation. We adopt the functional form of a modified Calzetti law (Noll2009), but it should not be interpreted as indicating any particular amount of dust attenuation. The parameterization can be expressed as:
where the free parameters are , , , , , , and . and are the amplitude and power-law slope of the UV continuum. It represents a distinct component, which could be either due to the AGN light or the host galaxy. denotes the form of the modified Calzetti law (Noll2009), where represents a generalized attenuation curve (not necessarily tied to physical dust properties), and is the power-law modification parameter, normalized at 5500 Å. For J1047+0739, we introduce an additional free parameter to model the Balmer jump as a discontinuity in the power-law at 3645 Å. is the blackbody spectrum at temperature , and is its amplitude. For the dust emission, we adopt the dust templates from Lyu2021. and are the amplitude and blackbody or gray body spectrum of the -th dust component.
The best-fit continua of the three local LRDs are shown in Figure D.2. The corresponding best-fit parameters and component luminosities are summarized in Table D.1. We do not include associated uncertainties given the descriptive nature of this parameterization. All three objects exhibit temperatures around 5000–6000 K and very negative values of . The extreme produces a very steep modification to the original blackbody. Such a negative strongly suppresses the UV-to-optical flux of the blackbody spectrum while leaving the IR spectrum largely unaffected. As discussed in §6.2, physical mechanisms, such as metal cooling and gas opacity, may contribute to this modification. In particular, the contribution of the UV power-law component to the optical continuum in J1047+0739 is non-negligible. At the H wavelength, the power-law component is only a factor of 2.3 fainter than the pseudo-blackbody. This is broadly consistent with measurements from its HST narrow-band image, which show the AGN luminosity to be 5.5 times that of the host galaxy. In the other two objects, the power-law components at optical wavelengths are orders of magnitude fainter than the pseudo-blackbody components.